You're four years older than your 18 year old sister. How many years ago were you three times her age? The correct answer is 16 years ago(you were 6 and she was 2)
Answer by Maths68(1474)

You're four years older than your 18 year old sister.
My age = 18 + 4
My age = 22
Now suppose x years ago I was three times of her age
x years ago means
my age was = 22-x
my sisters age was = 18-x
then I was three times of her age
22-x=3(18-x)
22-x=54-3x
-x+3x=54-22
2x=32
x=32/2
x=16

16 years ago I was three times of my sisters age.